OpenStudy (anonymous):

write an equation of the line that passes through the given points. (-3,1),(0,-8)

OpenStudy (anonymous):

Find the slope \[Slope = \frac{y2 - y1}{x2 - x1}\] and then use the slope and one of the given points and use the point-slope formula to get the equation of the line. \[y - y1 = m(x - x1)\]

OpenStudy (anonymous):

is that the answer?

OpenStudy (anonymous):

No but all you got to do is plug in numbers and you will get the answer.

OpenStudy (anonymous):

Ugh hat too much work lol

OpenStudy (anonymous):

Lol im sorry ;)

OpenStudy (anonymous):

ewww

OpenStudy (anonymous):

?

OpenStudy (anonymous):

?

OpenStudy (anonymous):

you said eww.. im confused lol

OpenStudy (anonymous):

oh my bad worng chat sorry